# Implicit Bias of SignGD and Adam on Multiclass Separable Data

### v1

@@ Abstract
In the optimization of overparameterized models, different gradient-based methods can achieve zero training error yet converge to distinctly different solutions inducing different generalization properties. While a decade of research on implicit optimization bias has illuminated this phenomenon in various settings, even the foundational case of linear classification with separable data still has important open questions. We resolve a fundamental gap by characterizing the implicit bias of both Adam and Sign Gradient Descent in multi-class cross-entropy minimization: we prove that their iterates converge to solutions that maximize the margin with respect to the classifier matrix’s max-norm and characterize the rate of convergence. We extend our results to general p-norm normalized steepest descent algorithms and to other multi-class losses.
Machine Learning, ICML
@@ 1 Introduction
Machine learning models are trained to minimize a surrogate loss function with the goal of finding model-weight configurations that generalize well. The loss function used for training is typically a proxy for the evaluation metric. The prototypical example is one-hot classification where the predominant choice, the cross-entropy (CE) loss, serves as a convex surrogate of the zero-one metric that measures correct class membership. Intuitively, a good surrogate loss is one where weight configurations minimizing the training loss also achieve strong generalization performance. However, modern machine learning models are overparameterized, leading to multiple weight configurations that achieve identical training loss but exhibit markedly different generalization properties (Zhang et al., 2017; Belkin et al., 2019).
This observation has motivated extensive research into the implicit bias (or implicit regularization) of gradient-based optimization. This theory investigates how optimizers select specific solutions from the infinite set of possible minimizers in overparameterized settings. The key insight is that gradient-based methods inherently prefer “simple” solutions according to optimizer-specific notions of simplicity. Understanding this preference requires analyzing not just loss convergence, but the geometric trajectory of parameter updates throughout training.
The prototypical example—which has rightfully earned its place as a “textbook” setting in the field—is the study of implicit optimization bias of gradient descent in linear classification. In this setting, embeddings of the training data are assumed fixed and only the classifier is learned, with overparameterization modeled through linear separability of the training data. (Soudry et al., 2018) demonstrated that while the GD parameters diverge in norm, they converge in direction to the maximum-margin classifier—the minimum $L_{2}$-norm solution that separates the data. This non-trivial result, despite (or perhaps because of) the simplicity of its setting, spurred numerous research directions extending to nonlinear models, alternative optimizers, different loss functions, and theoretical explanations of phenomena like benign overfitting (see Sec. 2).
Our paper contributes a fundamental result to this literature by establishing the implicit bias of Adam and its simpler variant, signed GD (SignGD), in the multiclass version of this textbook setting: minimization of CE loss on multiclass linearly separable data. Our work addresses several gaps between theory assumption and practice: the predominant focus on binary classification despite multiclass problems dominating real-world applications; the much theoretical emphasis on exponential loss despite practitioners’ overwhelming preference for cross-entropy loss; and the limited theoretical understanding of Adam’s implicit bias, which has only recently been studied in the binary case (Zhang et al., 2024) despite its widespread adoption in deep learning practice. Here, we provide a direct analysis of the multiclass setting by exploiting the nice properties of the softmax function, avoiding the traditional approach of reducing multiclass problems to binary ones (Wang & Scott, 2024).
@@
Contributions. Our contributions are as follows: For multiclass separable data trained with CE loss, we show that the iterates of SignGD converge to a solution that maximizes the margin defined with respect to (w.r.t.) the matrix max-norm, with a rate $\mathcal{O}(\frac{1}{t^{1/2}})$. We generalize this result to normalized steepest descent algorithms w.r.t. any entry-wise matrix $p$-norm. This directly extends Nacson et al. (2019); Sun et al. (2023)’s results to multiclass classification and to the widely-used CE loss. To achieve this, we construct a proxy for the loss and show that it closely traces its value and gradient. We also show the same machinery applies to other multiclass losses such as the exponential loss (Mukherjee & Schapire, 2010) and the PairLogLoss (Wang et al., 2021b).
Under the same setting, we prove that the iterates of Adam also maximize the margin w.r.t. the matrix max-norm, with a rate $\mathcal{O}(\frac{1}{t^{1/3}})$. This matches the convergence rate previously established for binary classification (Zhang et al., 2024), and remarkably, is independent of the number of classes $k$. This class-independence is non-trivial since naive reduction approaches that map a $k$-class problem in $d$ dimensions to a binary problem in $kd$ dimensions inevitably introduce $k$-dependent factors. Our key insight is to decompose the CE proxy function class-wise, enabling us to bound Adam’s first and second gradient moments separately for each class.This decomposition reveals how each component of the proxy function governs the dynamics of the weight vectors associated with its corresponding class. Finally, we experimentally verify our theory predictions for the algorithms considered. Specific to Adam, we numerically demonstrate that the implicit bias results are consistent with or without the (small) stability constant in the multiclass setting, and the solutions found by SignGD and Adam favor the max-norm margin over the 2-norm margin.
2 Related Works
Starting with GD, the foundational result by (Soudry et al., 2018) showed that gradient descent optimization of logistic loss on linearly separable data converges in direction to the $L_{2}$ max-margin classifier at a rate $O(1/\log(t))$. Contemporaneous work by (Ji & Telgarsky, 2019) generalized this by removing the data separability requirement. (Ji et al., 2020) later connected these findings to earlier work on regularization paths of logistic loss minimization (Rosset et al., 2003), which enabled extensions to other loss functions (e.g., those with polynomial tail decay). More recently, (Wu et al., 2024b) extends these results to the large step size regime with the same $O(1/\log(t))$ rate. The relatively slow convergence rate to the max-margin classifier motivated investigation into adaptive step-sizes. (Nacson et al., 2019) showed that normalized gradient descent (NGD) with decaying step-size $\eta_{t}=1/\sqrt{t}$ achieves $L_{2}$-margin convergence at rate $O(1/\sqrt{t})$. This rate was improved to $O(1/t)$ by (Ji & Telgarsky, 2021) using constant step-sizes, and further to $O(1/t^{2})$ through a specific momentum formulation (Ji et al., 2021). Besides linear classifications, implicit bias of GD has been studied for least squares (Gunasekar et al., 2017, 2018), homogeneous (Lyu & Li, 2019; Ji & Telgarsky, 2020) and non-homogeneous neural networks (Wu et al., 2024a), and matrix factorization (Gunasekar et al., 2017); see (Vardi, 2023) for a survey.
[…]
On the other hand, the implicit bias of adaptive algorithms such as Adagrad (Duchi et al., 2011) or Adam (Kingma & Ba, 2014) is less explored compared to GD. (Qian & Qian, 2019) studied the implict bias of Adagrad and showed its directional convergence to a solution characterized by a quadratic minimization problem. (Wang et al., 2021a, 2022) demonstrated the normalized iterates of Adam (with non-negligible stability constant) converge to a KKT point of a $L_{2}$-margin maximization problem for homogeneous neural networks. More recently and most relevant to our work, (Zhang et al., 2024) studied the implicit bias of Adam without the stability constant on binary linearly separable data. They showed that unlike GD, the Adam iterates converge to a solution that maximizes the margin with respect to the $L_{\infty}$-norm. This study excluding the stability constant is practically-relevant given the magnitude of the constant is typically very small (default $81e-$ 8 in PyTorch (Paszke et al., 2019)). This setting is also the focus of another closely-related recent study of the implicit bias of AdamW (Xie & Li, 2024), where the authors again establish that convergence aligns with the $L_{\infty}$ (rather than $L_{2}$) geometry. Our work extends these latter studies to the multiclass setting (see Remark 6.5 for technical comparisons).
All the above mentioned works focus solely on binary classification. The noticeable gap in analysis of multi class classification in most existing literature is recently emphasized by (Ravi et al., 2024) who extend the implicit bias result of (Soudry et al., 2018) to multiclass classification for losses with exponential tails, including cross-entropy, multiclass exponential, and PairLogLoss. Their approach leverages a framework introduced by (Wang & Scott, 2024), mapping multiclass analysis to binary cases. Our work directly addresses their open questions regarding the implicit bias of alternative gradient-based methods in multiclass settings by analyzing methods with adaptive step-sizes. Thanks to the adaptive step-sizes, our rates of convergence to the margin improve to polynomial dependence on $t$. Furthermore, our technical approach differs: rather than mapping to binary analysis, we work directly with multiclass losses, exploiting properties of the softmax function to produce elegant proofs that apply to all three losses studied by (Ravi et al., 2024). Our class-wise decomposition is crucial for analyzing Adam with the same convergence rate as the binary case, avoiding any extra factors that depend on the number of classes.
@@ 3 Preliminaries
For any integer $k$, $[k]$ denotes $\{1,\ldots,k\}$. Matrices, vectors, and scalars are denoted by $\bm{A}$, $\bm{a}$, and $a$ respectively. For matrix $\bm{A}$, we denote its $(i,j)$-th entry as $\bm{A}[i,j]$, and for vector $\bm{a}$, its $i$-th entry as $\bm{a}[i]$. We consider entry-wise matrix $p$-norms defined as $\|\bm{A}\|_{p}=(\sum_{i,j}|{\bm{A}}[i,j]|^{p})^{1/p}$. Central to our results are: the infinity norm, denoted as ${\left\|\bm{A}\right\|_{\max}}=\max_{i,j}|\bm{A}[i,j]|$ and called the max-norm, and the entry-wise 1-norm, denoted as ${\left\|{\bm{A}}\right\|_{\rm{sum}}}=\sum_{i,j}|\bm{A}[i,j]|$. For any other entry-wise $p$-norm with $p>1$, we write $\|{\bm{A}}\|$ (dropping subscripts) for simplicity. We denote by $\|\bm{A}\|_{*}$ the dual-norm with respect to the standard matrix inner product $\langle\bm{A},\bm{B}\rangle=\operatorname{tr}(\bm{A}^{\top}\bm{B})$. The entry-wise 1-norm is dual to the max-norm. For vectors, the max-norm is equivalent to the infinity norm, denoted as $\|\bm{a}\|_{\infty}$, while we denote the $\ell_{1}$ norm as $\|\bm{a}\|_{1}$. Let indicator $\delta_{ij}$ such that $\delta_{ij}=1$ if and only if $i=j$. Denote $\mathbb{S}:\mathbb{R}^{k}\rightarrow\triangle^{k-1}$ the softmax map of $k$-dimensional vectors to the probability simplex $\triangle^{k-1}$ such that for $\bm{a}\in\mathbb{R}^{k}$:
$\mathbb{S}(\bm{a})=\big{[}\frac{\exp({\bm{a}[c]})}{\sum_{c\in[k]}\exp(\bm{a}[c])}\big{]}_{c=1}^{k}\in\triangle^{k-1}.$
Let $\mathbb{S}_{c}({\bm{v}})$ denote the $c$-th entry of $\mathbb{S}({\bm{v}})$. Let $\mathbb{S}^{\prime}(\bm{a})=\operatorname{diag}(\mathbb{S}(\bm{a}))-\mathbb{S}(\bm{a})\mathbb{S}(\bm{a})^{\top}$ denote the softmax gradient, with $\operatorname{diag}(\cdot)$ a diagonal matrix. Finally, let $\{\bm{e}_{c}\}_{c=1}^{k}$ be the standard basis vectors of $\mathbb{R}^{k}$.
@@
Consider a multiclass classification problem with training data ${\bm{h}_{1},\ldots,\bm{h}_{n}}$ and labels ${y_{1},\ldots,y_{n}}$. Each datapoint $\bm{h}_{i}\in\mathbb{R}^{d}$ is a vector in a $d$-dimensional embedding space (denote data matrix ${\bm{H}}=[\bm{h}_{1},\ldots,\bm{h}_{n}]^{\top}\in\mathbb{R}^{n\times d}$), and each label $y_{i}\in[k]$ represents one of $k$ classes. We assume each class contains at least one datapoint. The classifier $f_{{\bm{W}}}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a linear model with weight matrix ${\bm{W}}\in\mathbb{R}^{k\times d}$. The model outputs logits $\bm{\ell}_{i}=f_{{\bm{W}}}(\bm{h}_{i})={\bm{W}}\bm{h}_{i}$ for $i\in[n]$, which are passed through the softmax map to produce class probabilities $\hat{p}(c|\bm{h}_{i})=\mathbb{S}_{c}(\bm{\ell}_{i})$. We train using empirical risk minimization (ERM): $\mathcal{L}_{\text{ERM}}({\bm{W}}):=-\frac{1}{n}\sum_{i\in[n]}\ell\left({{\bm{W}}\bm{h}_{i}};y_{i}\right)\,,$ where the loss function $\ell$ takes as input the logits of a datapoint and its label. The predominant choice in classification is the CE loss
$\mathcal{L}({\bm{W}}):=-\frac{1}{n}\sum\nolimits_{i\in[n]}\log\big{(}\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i})\big{)}$ (1)
$=\frac{1}{n}\sum\nolimits_{i\in[n]}\log\big{(}1+\sum\nolimits_{c\neq y_{i}}\exp(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i})\big{)}\,,$
@@
We focus our discussions on the CE loss due to its ubiquity in practice. However, our results hold for other multiclass losses such as the exponential (Mukherjee & Schapire, 2010) and the PairLogLoss (Wang et al., 2021b) (see App. F).

### v4

# Implicit Bias of Spectral Descent and Muon on Multiclass Separable Data

@@ Abstract
Different gradient-based methods for optimizing overparameterized models can all achieve zero training error yet converge to distinctly different solutions inducing different generalization properties. We provide the first complete characterization of implicit optimization bias for p-norm normalized steepest descent (NSD) and momentum steepest descent (NMD) algorithms in multi-class linear classification with cross-entropy loss. Our key theoretical contribution is proving that these algorithms converge to solutions maximizing the margin with respect to the classifier matrix’s p-norm, with established convergence rates. These results encompass important special cases including Spectral Descent and Muon, which we show converge to max-margin solutions with respect to the spectral norm. A key insight of our contribution is that the analysis of general entry-wise and Schatten p-norms can be reduced to the analysis of NSD/NMD with max-norm by exploiting a natural ordering property between all p-norms relative to the max-norm and its dual sum-norm. For the specific case of descent with respect to the max-norm, we further extend our analysis to include preconditioning, showing that Adam converges to the matrix’s max-norm solution. Our results demonstrate that the multi-class linear setting, which is inherently richer than the binary counterpart, provides the most transparent framework for studying implicit biases of matrix-parameter optimization algorithms.
@@ 1 Introduction
The ever-increasing training cost of large language models (LLMs) has demanded better optimizer designs with improved performance and efficiency [brown2020language, achiam2023gpt, guo2025deepseek]. The de facto standard optimizers for deep learning training are Adam and AdamW [kingma2014adam, loshchilov2017decoupled]. However, these algorithms that employ diagonal preconditioners to independently adjust the learning rate of each coordinate, may fail to capture their inter-dependencies and fully leverage the geometry of the loss landscape [zhang2024transformers]. This has spurred a series of research efforts on improving Adam or AdamW’s computational efficiency [grosse2016kronecker, gupta2018shampoo, shazeer2018adafactor, zhang2024adam], with LLM-training as the target application domain [jordan2024muon, vyas2024soap, pethick2025training, liu2025muon].
A noticeable work by jordan2024muon proposed the Muon optimizer, which was shown to have remarkable performances on NanoGPT benchmarks. More recently, it has been shown that Muon can be used for large-scale LLM training with the potential to replace AdamW as the standard choice [liu2025muon]. The key step in Muon is to orthogonalize the updates via the Newton-Schulz iteration [jordan2024muon, bernstein2024old]. More precisely, the update (denoted as $\bm{\Delta}$) is (approximately) replaced by the product of its singular-vector matrices $\bm{U}{\bm{V}}^{T}$ (where the (truncated) singular value decomposition (SVD) of $\bm{\Delta}$ is $\bm{\Delta}=\bm{U}\bm{\Sigma}{\bm{V}}^{T}$). Even though the benefits of orthogonalization are not fully understood, jordan2024muon pointed out that it could promote updates in directions of small magnitudes given the weight matrices are typically low-rank. Moreover, if the above SVD approximation is exact and gradient accumulations are turned off, then Muon becomes spectral descent [carlson2015preconditioned, bernstein2024old], which is the (normalized) steepest descent w.r.t the spectral norm [bernstein2024old]. As noted by bernstein2024old, spectral descent is also Shampoo (which won the AlgoPerf competition [shi2023distributed, dahl2023benchmarking]) without accumulations in preconditioners. Thus, Muon can be viewed as (approximate) Shampoo when both optimizers are without accumulations. In essence, we observe that one important ingredient of Muon and Shampoo (without accumulations) is the spectral-descent step,
$\bm{W}^{\dagger}={\bm{W}}-\eta\bm{U}{\bm{V}}^{T}\quad\text{where}\quad\nabla\mathcal{L}({\bm{W}})=\bm{U}\bm{\Sigma}{\bm{V}}^{T}.$
Theoretical investigations of spectral descent or Muon mainly focus on characterizing the convergence rates of the algorithm (e.g., the rate of decrease of the gradient norm in the non-convex setting [an2025asgo, li2025note, pethick2025training]). However, modern machine learning models are overparameterized, leading to multiple weight configurations that achieve identical training loss but exhibit markedly different generalization properties [zhang2017understanding, belkin2019reconciling]. The key insight is that gradient-based methods inherently prefer “simple” solutions according to optimizer-specific notions of simplicity. Understanding this implicit bias/regularization requires analyzing not just loss convergence, but the geometric trajectory of parameter updates throughout training. To this end, our work aims to address the fundamental question:
What is the implicit bias of spectral descent (and its momentum variants) in linear multiclass classification with separable data and cross-entropy loss?
Table 1: Summary of margin convergence rates for NSD and NMD algorithms of different norm constraints for linear multiclass separable data with the CE loss. The (truncated) SVDs of the gradient and momentum are denoted as $\bm{\nabla}=\bm{U}\bm{\Sigma}{\bm{V}}^{T}$ and $\bm{M}=\tilde{\bm{U}}\tilde{\bm{\Sigma}}\tilde{{\bm{V}}}^{T}$ respectively.
Method Norm Constraint Update $\bm{\Delta}$ Reference Rate 2
NGD Unit $\lVert\cdot\rVert_{2}$-ball $\frac{\bm{\nabla}}{\lVert\bm{\nabla}\rVert_{2}}$ hazan2015beyond -
NMD-GD $\frac{\bm{M}}{\lVert\bm{M}\rVert_{2}}$ cutkosky2020momentum -
SignGD Unit $\left\|\cdot\right\|_{\max}$}-ball $\texttt{sign}(\bm{\nabla})$ bernstein2018signsgd -
Signum $\texttt{sign}(\bm{M})$ bernstein2018signsgd -
Spectral-GD Unit $\left|\left|\left|\cdot\right|\right|\right|_{\infty}$-ball $\bm{U}{\bm{V}}^{T}$ bernstein2024old $O(\frac{\log t+n}{t^{1/2}})$
Muon 1     $\tilde{\bm{U}}\tilde{{\bm{V}}}^{T}$   jordan2024muon $O(\frac{d\log t+dn}{t^{1/2}})$
  • 1
    We consider EMA-style momentum of the form (5).
  • 2
    NGD and SignGD rates are the same as Spectral-GD; Signum and NMD-GD rates are the same as Muon.
The multiclass setting where the parameter is a matrix, is a natural place to study the class of spectral-descent algorithms, and provides an inherently richer setting. Our work captures this richness by establishing convergence with respect to not only entry-wise matrix norms, but also matrix Schatten norms. Hence, while the focus is on spectral descent and Muon, the analysis establishes implicit bias rates for a wide family of algorithms (Table 1), and we state the results in the most general form from the perspective of steepest descent with (unit) norm-ball constraints. Our contributions are as follows:
  1. 1.
@@
    For multiclass separable data trained with the cross-entropy (CE) loss, we show that the iterates of normalized steepest descent (NSD) defined with respect to (w.r.t.) any matrix entry-wise or Schatten norms converge to a solution that maximizes the margin defined w.r.t. the same norm, with a rate $\mathcal{O}(\frac{1}{t^{1/2}})$. This includes sign descent (entry-wise max-norm) [bernstein2018signsgd], normalized gradient descent (entry-wise 2-norm) [hazan2015beyond], and spectral descent (Schatten $\infty$-norm) [bernstein2024old] as special cases. To achieve this, we introduce a unified analysis framework that relates entry-wise and Schatten p-norms to the entry-wise max-norm, and construct a proxy function for the loss that closely traces both its value and gradient. We also show the same machinery applies to other multiclass losses such as the exponential loss [mukherjee2010theory] and the PairLogLoss [wang2021rank4class].
  2. 2.
    Under the same setting, we utilize the same framework and proxy function to show that a $\mathcal{O}(\frac{1}{t^{1/2}})$ margin convergence rate also holds for normalized momentum steepest descent (NMD). This includes the following algorithms in analogy to the ones above: sign momentum descent [bernstein2018signsgd], normalized momentum gradient descent [cutkosky2020momentum], and Muon [jordan2024muon]. The key step of the analysis is to use the proxy function to bound the sum-norm difference between the gradient and the momentum (i.e., the exponential moving averages (EMA) of the gradient), which translates to a bound on the dual norm through the fundamental norm-relationships used in the study of NSD. The margin convergence rates of various algorithms are summarized in Table 1. Furthermore, we extend the analysis to Adam (without the stability constant) and show its iterates maximize the margin w.r.t. the matrix max-norm (proof details and numerical validations in App. G).
  3. 3.
    We experimentally verify our theoretical predictions across all considered algorithms. First, for sign descent (SignGD) and Signum, we demonstrate that solutions favor the max-norm margin over the 2-norm margin—the opposite behavior to normalized gradient descent (NGD) and normalized momentum gradient descent (NMD-GD). Moreover, we show that both spectral descent (Spectral-GD) and Muon favor the spectral-norm margin over the other norms. We further extend the experiments to the non-linear setting with a two-layer neural network. We observe the (unnormalized) spectral-norm margin of Spectral-GD and Muon grow faster than other algorithms. Hence, the norm-preference trend in the linear setting can also exhibit in the non-linear setting.
@@ 2 Preliminaries
[…]
Let $\mathbb{S}:\mathbb{R}^{k}\rightarrow\triangle^{k-1}$ the softmax map of $k$-dimensional vectors to the probability simplex $\triangle^{k-1}$ such that for any $\bm{a}\in\mathbb{R}^{k}$, it holds that $\mathbb{S}(\bm{a})=\big[\frac{\exp({\bm{a}[c]})}{\sum_{c\in[k]}\exp(\bm{a}[c])}\big]_{c=1}^{k}\in\triangle^{k-1}$. Let $\mathbb{S}_{c}({\bm{v}})$ denote the $c$-th entry of $\mathbb{S}({\bm{v}})$. Let $\mathbb{S}^{\prime}(\bm{a})=\operatorname{diag}(\mathbb{S}(\bm{a}))-\mathbb{S}(\bm{a})\mathbb{S}(\bm{a})^{\top}$ denote the softmax gradient, with $\operatorname{diag}(\cdot)$ a diagonal matrix. Finally, let $\{\bm{e}_{c}\}_{c=1}^{k}$ be the standard basis vectors of $\mathbb{R}^{k}$, and indicator $\delta_{ij}$ be such that $\delta_{ij}=1$ if and only if $i=j$. For any integer $k$, $[k]$ denotes $\{1,\ldots,k\}$.
@@
Consider a multiclass classification problem with training data $\bm{h}_{1},\ldots,\bm{h}_{n}$} and labels $y_{1},\ldots,y_{n}$}. Each datapoint $\bm{h}_{i}\in\mathbb{R}^{d}$ is a vector in a $d$-dimensional embedding space (denote data matrix ${\bm{H}}=[\bm{h}_{1},\ldots,\bm{h}_{n}]^{\top}\in\mathbb{R}^{n\times d}$), and each label $y_{i}\in[k]$ represents one of $k$ classes. We assume each class contains at least one datapoint. The classifier $f_{{\bm{W}}}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a linear model with weight matrix ${\bm{W}}\in\mathbb{R}^{k\times d}$. The model outputs logits $\bm{\ell}_{i}=f_{{\bm{W}}}(\bm{h}_{i})={\bm{W}}\bm{h}_{i}$ for $i\in[n]$, which are passed through the softmax map to produce class probabilities $\hat{p}(c|\bm{h}_{i})=\mathbb{S}_{c}(\bm{\ell}_{i})$.
We train using empirical risk minimization (ERM): $\mathcal{L}_{\text{ERM}}({\bm{W}}):=-\frac{1}{n}\sum_{i\in[n]}\ell\left({{\bm{W}}\bm{h}_{i}};y_{i}\right)\,$, where the loss function $\ell$ takes as input the logits of a datapoint and its label. The predominant choice in classification is the CE loss
$\mathcal{L}({\bm{W}}):$ $=-\frac{1}{n}\sum\nolimits_{i\in[n]}\log\big(\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i})\big).$ (1)
@@
We focus our discussions on the CE loss due to its ubiquity in practice. However, our results hold for other multiclass losses such as the exponential [mukherjee2010theory] and the PairLogLoss [wang2021rank4class] (see App. F). Define the maximum margin of the dataset w.r.t. any entry-wise or Schatten p-norm $\|\cdot\|$ as
$\gamma:=\max\nolimits_{\|{\bm{W}}\|\leq 1}\,\min\nolimits_{{i\in[n],\,c\neq{y_{i}}}}\,\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}{\bm{W}}\bm{h}_{i}\,.$ (2)
Optimization Methods
We study iterative algorithms that update the weight matrix by
${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t}.$
For the NSD family [boyd2004convex], the update direction [^1] For $p\in(1,\infty)$, the norms $\|\cdot\|_{p}$ and $\left|\left|\left|\cdot\right|\right|\right|_{p}$ are strictly convex, thus there is a unique maximizer defining the update in Eqn. (3). For $p=1,\infty$ the maximizer is not necessarily unique and our results hold for any choice of $\bm{\Delta}_{t}$ in the set of maximizers; see e.g. ziketak1988characterization. w.r.t. the norm $\lVert\cdot\rVert$ is
$\bm{\Delta}_{t}:=\arg\max\nolimits_{\|\bm{\Delta}\|\leq 1}\langle\bm{\nabla}_{t},\bm{\Delta}\rangle\,.$ (3)
Note that this reduces to SignGD, Coordinate Descent (e.g., nutini2015coordinate), or NGD when the max-norm (i.e. $\lVert\cdot\rVert_{\infty}$), the entry-wise $1$-norm (i.e. $\left\|\cdot\right\|_{\rm{sum}}$}), or the Frobenius Euclidean-norm (i.e. $\lVert\cdot\rVert_{2}$) is used, respectively. Concretely, the update directions for SignGD and NGD are:
$\text{SignGD:}\quad\bm{\Delta}_{t}=\texttt{sign}(\bm{\nabla}_{t}),\quad\text{and}\quad\text{NGD:}\quad\bm{\Delta}_{t}=\bm{\nabla}_{t}/\lVert\bm{\nabla}_{t}\rVert_{2},$
where the $\texttt{sign}(\cdot)$ and division $\frac{\cdot}{\cdot}$ operations are applied entry-wise. In the special case of spectral norm (i.e. $\left|\left|\left|\cdot\right|\right|\right|_{\infty}$), this becomes the Spectral-GD, for which $\bm{\Delta}_{t}=\bm{U}_{t}{\bm{V}}_{t}^{T}$, where $\bm{U}_{t}$ and ${\bm{V}}_{t}$ are the left/right singular matrices of $\bm{\nabla}_{t}$ respectively (i.e., $\bm{\nabla}_{t}=\bm{U}_{t}\bm{\Sigma}_{t}{\bm{V}}_{t}^{T}$ with singular values in $\bm{\Sigma}_{t}>0$ arranged in non-increasing order). Finally, note that the Schatten $2$-norm case reduces to NGD (as $\left|\left|\left|\cdot\right|\right|\right|_{2}=\lVert\cdot\rVert_{2}$).
We also consider the NMD family with the following update direction w.r.t. the norm $\lVert\cdot\rVert$
$\bm{\Delta}_{t}:=\arg\max\nolimits_{\lVert\bm{\Delta}\rVert\leq 1}\langle\bm{M}_{t},\bm{\Delta}\rangle\,$, (4)
where the momentum $\bm{M}_{t}$ is computed as the EMA of the gradient given by
$\mathbf{M}_{t}$ $=\beta_{1}\mathbf{M}_{t-1}+(1-\beta_{1})\bm{\nabla}_{t}.$ (5)
This form of momentum is also known as the heavy-ball or the SGDM-style momentum [polyak1964some, ghadimi2015global, liu2020improved]. Thus, an NMD algorithm chooses the update direction (among all feasible directions in the unit $\lVert\cdot\rVert$-ball) that best aligns with the momentum instead of the gradient direction (as chosen by an NSD algorithm). Similar to above, when the max-norm and the Frobenius-norm are used, the resulting Signum and NMD-GD update directions are:
$\text{Signum:}\quad\bm{\Delta}_{t}=\texttt{sign}(\bm{M}_{t}),\quad\text{and}\quad\text{NMD-MD:}\quad\bm{\Delta}_{t}={\bm{M}_{t}}\big/{\lVert\bm{M}_{t}\rVert_{2}}.$
When spectral norm is used in (4), this becomes Muon [^2] The implementation in jordan2024muon uses Nesterov-type momentum: Newton-Schulz iteration applied to $\beta_{1}\bm{M}_{t}+\bm{\nabla}_{t}$ instead of $\beta_{1}\bm{M}_{t-1}+\bm{\nabla}_{t}$ [liu2025muon]. for which the SVD is on $\bm{M}_{t}$ (i.e. $\bm{M}_{t}=\tilde{\bm{U}}_{t}\tilde{\bm{\Sigma}}_{t}\tilde{{\bm{V}}}_{t}^{T}$) and the update direction is $\bm{\Delta}_{t}=\tilde{\bm{U}}_{t}\tilde{{\bm{V}}}_{t}^{T}$. Note that Muon reduces to Spectral-GD when the momentum parameter $\beta_{1}$ is set to 0. Similar reductions also hold for Signum (to SignGD) and NMD-GD (to NGD) as well.
Assumptions
Establishing the implicit bias of the above mentioned gradient-based optimization algorithms requires the following assumptions. First, we assume data are linearly separable, ensuring the margin $\gamma$ is strictly positive, an assumption routinely used in previous works [soudry2018implicit, ravi2024implicit, gunasekar2018characterizing, nacson2019convergence, wu2024implicit].
Assumption 1.
In this work, we consider learning rate schedule $\eta_{t}=\Theta(\frac{1}{t^{a}})$, where $a\in(0,1]$. Such schedules have been studied in the convergence and implicit bias of various optimization algorithms (e.g., bottou2018optimization, nacson2019convergence, and sun2023unified) including Adam [zhang2024implicit].
Assumption 2.
Assumption 3 can be satisfied by the above learning rate for a sufficiently large $t$ as shown in zhang2024implicit. It is used in our analysis of NMD and Adam.
Assumption 3.
Finally, we assume that the $1$-norm of the data is bounded. Similar assumptions were used in ji2019implicit, nacson2019convergence, wu2024implicit, and zhang2024implicit.
Assumption 4.
3 A Unified Framework with a Proxy Function
Analyzing margin convergence begins with studying loss convergence through second-order Taylor expansion of the CE loss (recall that ${\mathbb{S}}^{\prime}({\bm{v}})=\operatorname{diag}({\bm{v}})-{\bm{v}}{\bm{v}}^{\top}$):
$\mathcal{L}({\bm{W}}+\bm{\Delta})=\mathcal{L}({\bm{W}})+\langle\nabla\mathcal{L}({\bm{W}}),\bm{\Delta}\rangle+\frac{1}{2n}\sum\nolimits_{i\in[n]}\bm{h}_{i}^{\top}\bm{\Delta}^{\top}{\mathbb{S}}^{\prime}({\bm{W}}\bm{h}_{i})\bm{\Delta}\bm{h}_{i}+o(\|\bm{\Delta}\|_{F}^{3}),$ (6)
To bound the loss at ${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t}$, we must bound both the first-order and second-order terms in (6). For NSD updates in Eq. (3), the first term evaluates to $\eta_{t}\|\nabla\mathcal{L}({\bm{W}})\|_{*}$ (recall that $\|\cdot\|_{*}$ is the dual norm). This leads to two key tasks: (1) Lower-bounding the dual gradient norm; (2) Upper-bounding the second-order term.
For the proof to proceed, these bounds should satisfy two desiderata: (1) They are expressible as the same function of $\bm{W}$}, call it ${\mathcal{G}}({\bm{W}})$, up to constants. (2) The function ${\mathcal{G}}({\bm{W}})$ is a good proxy for the loss for small values of the latter. The former helps with combining the terms, while the latter helps with demonstrating descent. Next, we obtain these key bounds for the CE loss by determining the appropriate proxy ${\mathcal{G}}({\bm{W}})$.
Besides the need for a proxy ${\mathcal{G}}({\bm{W}})$, we use the following facts about the sum-norm dominating any entry-wise/Schatten p-norm. Concretely, for any matrix $\bm{A}$ and any $p\geq 1$: